# Ordinary primes for $\operatorname{GL}_{2}$-type abelian varieties and weight $2$ modular forms

### v1

@@ 2.2. Splitting of primes
Let $K/E$ be a finite extension of number fields and let $\widetilde{K}/E$ be the Galois closure of $K/E$. Let $G=\operatorname{Gal}(\widetilde{K}/E)$, $H=\operatorname{Gal}(\widetilde{K}/K)$, and $\Sigma=\{\sigma:K\rightarrow\widetilde{K}\;\big{|}\;\sigma|_{E}=\operatorname{id}_{E}\}$. Let $G$ (and hence $H$) act on $\Sigma$ via pre-composition. Write $n=[K:E]$. Then,
@@
    Let $\sigma_{n}\in\Sigma$ be the unique element such that $\sigma_{n}=\operatorname{id}_{K}$. Then, $H=\operatorname{Stab}_{G}(\{\sigma_{n}\})$ and the action of $H$ on $\Sigma\backslash\{\sigma_{n}\}$ induces an embedding $H\hookrightarrow\mathrm{Sym}(\Sigma\backslash\{\sigma_{n}\})$.
@@
For the third part, it suffices to show that $H=\operatorname{Stab}_{G}(\{\sigma_{n}\})$ so that it follows from the faithful action that $H\hookrightarrow\mathrm{Sym}(\Sigma\backslash\{\sigma_{n}\})$. Note that any element $\sigma\in\Sigma$ is uniquely determined by $\sigma(\alpha)$. Hence, $g\in\operatorname{Stab}_{G}(\{\sigma_{n}\})$ is equivalent to $g(\sigma_{n}(\alpha))=\sigma_{n}(\alpha)$, i.e., $g(\alpha)=\alpha$. As $K=E(\alpha)$, $g(\alpha)=\alpha$ is then equivalent to $g|_{K}=\operatorname{id}_{K}$, which is equivalent to $g\in H$. Hence, $H=\operatorname{Stab}_{G}(\{\sigma_{n}\})$.
@@
Let $L/K/E$ be a tower of finite extensions of number fields such that $L/E$ is a Galois extension. Let $G=\operatorname{Gal}(L/E)$ and $H=\operatorname{Gal}(L/K)$. Let $p$ be a prime of $E$ that is unramified in $L/E$ and let $\mathfrak{p}$ be a prime of $L$ lying above $p$. Let $G_{\mathfrak{p}}=\{\sigma\in G:\sigma(\mathfrak{p})=\mathfrak{p}\}$ and define an action of $G_{\mathfrak{p}}$ on $H\backslash G$ by $H\sigma\cdot g=H\sigma g$. Then,
@@
    there is a one-to-one correspondence between orbits of $H\backslash G$ under the action of $G_{\mathfrak{p}}$ and primes of $K$ that lie above $p$;
@@
    the inertia degree of the prime of $K$ corresponding to an orbit of $H\backslash G$ is the size of the orbit.
@@ 2.3. Extensions of degree $q$ or $2q$
Case 2: $v_{q}(|G|)=2$. By Sylow theorems, there exists a Sylow $q$-subgroup $I$ in $G$ of order $q^{2}$. If $q=2$, then $I$ can only be conjugate to $\langle(1,2,3,4)\rangle$, $\langle(1,2),(3,4)\rangle$, or $\langle(1,2)(3,4),(1,3)(2,4)\rangle$ in $S_{4}$. In particular, $I$ always contains an element which is a product of two disjoint cycles of length $q$. Now, suppose that $q\neq 2$. Note that $\langle(1,2,\ldots,q),(q+1,\ldots,2q)\rangle$ is a Sylow $q$-subgroup in $S_{2q}$ of order $q^{2}$, so $I$ is conjugate to $\langle(1,2,\ldots,q),(q+1,\ldots,2q)\rangle$ in $S_{2q}$ by Sylow theorems. Hence, $I$ is generated by two single cycles of length $q$ with disjoint elements. In particular, $I$ contains an element which is a product of two disjoint cycles of length $q$.
@@
[…]
@@ 3.1. Galois representations of abelian varieties
In this subsection, we will introduce the general theory of $\ell$-adic and $\lambda$-adic Galois representations of abelian varieties with extra endomorphisms. We will then specialize to the case of “$\operatorname{GL}_{2}$-type” abelian varieties, prove a result on the images of their Galois representations, and give a criterion, distinct from the definition, to identify their ordinary primes.
@@
which we call the $\ell$-adic Galois representation of $A$ [^1] Due to the symplectic pairing on $V_{\ell}(A)$ coming from the polarization of $A$, the image of $\rho_{A,\ell}$ actually lies in $\operatorname{GSp}_{2g}(\mathbb{Q}_{\ell})$. This Galois representation satisfies that for all primes $v$ of $F$ with $v\nmid N_{A}\ell$, $\rho_{A,\ell}$ is unramified at $v$, and the characteristic polynomial
@@
    $A/\overline{F}$ is of type I or II in Albert’s classification and $\operatorname{End}_{F}(A)\otimes\mathbb{Q}=\operatorname{End}_{\overline{F}}(A)\otimes\mathbb{Q}$.
@@
Let $A/F$ be an abelian variety of dimension $g$ over a number field $F$. Assume that $A$ is absolutely simple, $\operatorname{End}_{\overline{F}}(A)\otimes\mathbb{Q}$ contains a number field $K$ of degree $g$, [^4] Note that this is weaker than the assumption $K\hookrightarrow\operatorname{End}_{F}(A)\otimes\mathbb{Q}$ in Theorem 1.3. and the center of $\operatorname{End}_{\overline{F}}(A)\otimes\mathbb{Q}$ is a totally real field $K^{\prime}$. Then, there exists a finite extension $F^{\prime}/F$ such that for all sufficiently large prime $\ell$, we have, up to conjugation, that
@@
Let $F^{\prime}/F$ be a finite extension of number fields such that  [^5] The existence of this finite extension can be deduced from [Sil92, Theorem 4.1] and [SZ95, Lemma 2.7 and Remark 3.1] (see also [Ser00]).
@@
    $A$ is principally polarized over $F^{\prime}$;
  2. (2)
    $D:=\operatorname{End}_{F^{\prime}}(A)\otimes\mathbb{Q}=\operatorname{End}_{\overline{F}}(A)\otimes\mathbb{Q}$;
@@
By the assumptions and (2), $A/F^{\prime}$ is simple and of $\operatorname{GL}_{2}$-type, and by [Wu20, Proposition 1.7], $A/F^{\prime}$ is not of type III. As the center $K^{\prime}$ of $D$ is assumed to be totally real, by [Wu20, Proposition 1.7], we have either $D=K^{\prime}(=K)$ or $D$ is a division quaternion algebra over $K^{\prime}$ with $[K^{\prime}:\mathbb{Q}]=g/2$. Putting all things together, it is then easy to check that $A/F^{\prime}$ is of class $\mathcal{A}$, so by [BGK06, Theorem 6.16 (6.18)], we have $\operatorname{SL}_{2}(\mathcal{O}_{K^{\prime}}\otimes\mathbb{Z}_{\ell})\subseteq\rho_{A,\ell}(\operatorname{Gal}(\overline{F^{\prime}}/F^{\prime}))$ for sufficiently large prime $\ell$. On the other hand, by the assumptions, $A/F^{\prime}$ is (geometrically) of the first kind. As the center $K^{\prime}$ of $D$ is totally real, by Proposition 3.4 (and gathering all primes $\lambda^{\prime}$ of $K^{\prime}$ with $\lambda^{\prime}|\ell$), we have $\operatorname{det}\rho_{A,\ell}|_{\operatorname{Gal}(\overline{F}/F^{\prime})}=\chi_{\ell}$ for all rational primes $\ell$.
@@
Let $A$ be an abelian variety of dimension $g$ and conduct $N_{A}$ defined over a number field $F$. We say that a prime $v$ of $F$ is an ordinary prime of $A/F$ if $v\nmid N_{A}$ and $p\nmid a_{v,g}(A)$, where $p$ is the rational prime lying below $v$ and $a_{v,g}(A)$ is the coefficient of the term $X^{g}$ of $P_{A,v}(X)$ defined by Equation 1.
@@
Since #$\#\{\sigma:K\rightarrow\mathbb{C}\}=g$, it is easy to see that the coefficient $a_{v,g}(A)$ of the term $X^{g}$ is of the form
@@ 3.2. Galois representations of modular forms
such that for every rational primes $p\nmid\ell N$, $\rho_{f,\ell}$ is unramified at $p$ and
@@
By Ramanujan’s conjecture, it is known that $|\iota(a_{p}(f))|_{\mathbb{C}}\leq 2p^{\frac{k-1}{2}}$ for all embeddings $\iota:K\hookrightarrow\mathbb{C}$.
@@
Since #$\#\{\sigma:K_{f}\rightarrow\mathbb{C}\}=[K_{f}:\mathbb{Q}]=:g$, it is easy to see that the coefficient $a_{p,g}(A_{f})$ of the term $X^{g}$ is of the form
@@
where we write $[K:\mathbb{Q}]=g$. By the assumptions on $v$ and $p$, we have $\mathrm{N}(v)=p$ and $\mathrm{N}(\lambda)=p^{g}$. Hence, $p^{g}|\mathrm{N}((a_{v}))$ with $\mathrm{N}((a_{v}))\leq 2^{g}p^{g/2}$. As $p\geq 5$, this implies that $\mathrm{N}((a_{v}))=0$ so $a_{v}=0$. ∎
@@
By Lemma 4.1, if $v\in S_{1}$, then $a_{v}=0$ or $v\in S^{\mathrm{ord}}$. Since $\{v\in\Sigma_{F}:v\notin S\text{ and }a_{v}=0\}$ is assumed to have density 00, $S_{1}$ thus has the same density as $S_{1}\cap\{v\in\Sigma_{F}:v\notin S\text{ and }a_{v}\neq 0\}\subseteq S^{\mathrm{ord}}$, so it suffices to prove that $S_{1}$ has positive density.
@@
#$\#\{v\in\Sigma_{F}:\mathrm{N}(v)\leq X\}\sim\frac{X}{\log X}\sim\#\{p:p\leq X\}.$
@@
The proof of Theorem 1.6 follows a similar approach to that of Theorem 1.7. We first prove a lemma showing that under the assumption on the rational prime $p$, almost every prime $v$ of $F$ above $p$ is either an ordinary prime or belongs to a special set. In the case of Theorem 1.7, the special set is $\{v\in\Sigma_{F}:v\notin S\text{ and }a_{v}=0\}$, which has density 0 by assumptions. For Theorem 1.6, the special set is S(c) (see Equation 4 for its precise definition) or, more accurately, a finite union of such sets. It is much trickier to show that the density of $S(c)$ is 00. Therefore, we first proceed with the proof assuming that $S(c)$ has density 00, and then, in Section 5.2, we discuss its density properly.
@@
where we write $[K:\mathbb{Q}]=g$. By the assumptions on $v$ and $p$, we have $\mathrm{N}(v)=p$ and $\mathrm{N}(\lambda)=p^{g/2}$ ($g$ is necessarily even in this case). Hence, $p^{g/2}|\mathrm{N}((a_{v}))$ with $\mathrm{N}((a_{v}))\leq 2^{g}p^{g/2}$, implying that $\mathrm{N}_{K/\mathbb{Q}}(a_{v})^{2}=c\mathrm{N}(v)^{g}$ for some $c\in\mathbb{Z}$ with $0\leq c\leq 2^{2g}$. ∎
@@
Suppose for now that $S(c)$ has density 00, which will be proved in Corollary 5.7. We will show under this assumption that Theorem 1.6 holds. Let $T_{2}$ be the set of rational primes $p$ such that
@@
By Lemma 5.1, if $v\in S_{2}$, then $v\in S^{\mathrm{ord}}$ or $v\in S(c)$ for some $c\in\mathbb{Z}$ with $0\leq c\leq 2^{2g}$. As $S(c)$ and hence $\bigcup_{0\leq c\leq 2^{2g}}S(c)$ have density 00, it follows that $S_{2}$ has the same density as $S_{2}\cap S^{\mathrm{ord}}\subseteq S^{\mathrm{ord}}$, so it suffices to prove that $S_{2}$ has positive density. Now, the same argument as before gives that
@@ 5.2. Density of $S(c)$
The main goal of this subsection is to show that $S(c)$ has density 00 for each $c\in\mathbb{Z}$ with $0\leq c\leq 2^{2g}$, so that Theorem 1.6 holds by Section 5.1. To prove this, we need to use the surjective property of the family of Galois representations $(\rho_{\ell})_{\ell\in T}$.
Fix a rational prime $\ell$ that splits completely in $K$ and fix an embedding $\iota_{\ell}:\overline{\mathbb{Q}}\hookrightarrow\overline{\mathbb{Q}_{\ell}}$. Write $[K^{\prime}:\mathbb{Q}]=g^{\prime}$ and let $\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{g^{\prime}}:K^{\prime}\hookrightarrow\overline{\mathbb{Q}}$ be all the embeddings of $K^{\prime}$ into $\overline{\mathbb{Q}}$. As $\ell$ splits completely in $K$ and hence in $K^{\prime}$, by Lemma 2.2, $\{\lambda^{\prime}_{i}:=\lambda^{\prime}_{\sigma^{\prime}_{i}}\mid 1\leq i\leq g^{\prime}\}$ is an enumeration of all the places $\lambda^{\prime}|\ell$ of $K^{\prime}$. In particular, we may simply make the identification $K^{\prime}_{\lambda^{\prime}_{i}}=\widehat{\iota_{\ell}\sigma^{\prime}_{i}(K^{\prime})}=\mathbb{Q}_{\ell}\subseteq\overline{\mathbb{Q}_{\ell}}$ so that the natural map $K^{\prime}\rightarrow K^{\prime}_{\lambda^{\prime}_{i}}$ is given by $\alpha\mapsto\iota_{\ell}\sigma^{\prime}_{i}(\alpha)$. In this way,
@@
Write $[K:K^{\prime}]=d$. Since $\ell$ splits completely in $K$, each $\lambda^{\prime}_{i}$ also splits completely in $K$. In particular, for every prime $\lambda|\lambda^{\prime}_{i}$ of $K$, we have $\mathcal{O}_{K^{\prime}_{\lambda^{\prime}_{i}}}=\mathcal{O}_{K_{\lambda}}$. Let us now enumerate the primes $\lambda|\ell$ of $K$ by $(\lambda_{ij})_{i=1,j=1}^{g^{\prime},\;d}$ such that $\prod_{j=1}^{d}\lambda_{ij}=\lambda^{\prime}_{i}$ and let $\sigma_{ij}:K\hookrightarrow\overline{\mathbb{Q}}$ be the embedding corresponding to the prime $\lambda_{ij}$ by Lemma 2.2. We also make the identification $K_{\lambda_{ij}}=\widehat{\iota_{\ell}\sigma_{ij}(K)}=\mathbb{Q}_{\ell}\subseteq\overline{\mathbb{Q}_{\ell}}$ so that similarly
@@
[…]
Now, since $[F^{\prime}:F]$ is finite, there exists $n\in\mathbb{Z}^{+}$ such that $\tau^{n}\in\operatorname{Gal}(\overline{F}/F^{\prime})$ for all $\tau\in\operatorname{Gal}(\overline{F}/F)$. [^8] By considering all the cosets $\{\tau^{i}\operatorname{Gal}(\overline{F}/F^{\prime})\}_{i\in\mathbb{Z}}$, one can deduce that $\tau^{i}\in\operatorname{Gal}(\overline{F}/F^{\prime})$ for some $1\leq i\leq[F^{\prime}:F]$, so $n$ can be chosen as $[F^{\prime}:F]!$. In particular, this means that
@@
[…]
@@
[…]
@@
[…]
@@
We will now prove that the latter expression goes to 00 as $\ell\to\infty$, from which it follows that $S(c)$ has density 00.
@@
#$\#\left\{N\in\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z}):\operatorname{det}N=d,\operatorname{tr}N=t\right\}=\ell^{2}+\ell\cdot\bigg{(}\frac{t^{2}-4d}{\ell}\bigg{)},$
@@
$|C_{\ell}(g^{\prime},d,n,c_{0})|\leq n^{g^{\prime}d}\sum_{c\in\mu_{n}(\mathbb{F}_{\ell})}|D_{\ell}(g^{\prime},d,n,c{\hskip 1.0pt}\overline{c_{0}})|,$ (8)

### v2

@@ 2.2. Splitting of primes
Let $K/E$ be a finite extension of number fields and let $\widetilde{K}/E$ be the Galois closure of $K/E$. Let $G=\operatorname{Gal}(\widetilde{K}/E)$, $H=\operatorname{Gal}(\widetilde{K}/K)$, and $\Sigma=\{\sigma:K\rightarrow\widetilde{K}\;\big|\;\sigma|_{E}=\operatorname{id}_{E}\}$. Let $G$ (and hence $H$) act on $\Sigma$ via composition. Write $n=[K:E]$. Then,
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 2.3. Extensions of degree $q$ or $2q$
Case 2: $v_{q}(|G|)=2$. By Sylow theorems, there exists a Sylow $q$-subgroup II in $G$ of order $q^{2}$. If $q=2$, then $I$ can only be conjugate to $\langle(1,2,3,4)\rangle$, $\langle(1,2),(3,4)\rangle$, or $\langle(1,2)(3,4),(1,3)(2,4)\rangle$ in $S_{4}$. In particular, $I$ always contains an element which is a product of two disjoint cycles of length $q$. Now, suppose that $q\neq 2$. Note that $\langle(1,2,\ldots,q),(q+1,\ldots,2q)\rangle$ is a Sylow $q$-subgroup in $S_{2q}$ of order $q^{2}$, so $I$ is conjugate to $\langle(1,2,\ldots,q),(q+1,\ldots,2q)\rangle$ in $S_{2q}$ by Sylow theorems. Hence, $I$ is generated by two single cycles of length $q$ with disjoint elements. In particular, $I$ contains an element which is a product of two disjoint cycles of length $q$.
@@
Let $\widetilde{K}/E$ be the Galois closure of $K/E$. Write $n=[K:E]$, $G=\operatorname{Gal}(\widetilde{K}/E)$, and view $G\subseteq S_{n}$ by Lemma 2.3. Suppose first that $[K:E]=q$. By Lemma 2.6, it suffices to show that $G$ contains a cycle of length $q$. Note that $q$ divides $|G|$ since $G$ is a transitive subgroup of $S_{q}$. As $q$ is a prime, $G$ contains an element of order $q$, which is necessarily a cycle of length $q$ since $G\subseteq S_{q}$. Now, suppose that $[K:E]=2q$. By Lemma 2.6, it suffices to show that $G$ contains an element that is a product of two disjoint cycles of length $q$, and indeed this simply follows from Lemma 2.8. ∎
@@ 3.1. Galois representations of abelian varieties
In this subsection, we will introduce the general theory of $\ell$-adic and $\lambda$-adic Galois representations of abelian varieties with extra endomorphisms. We will then specialize to the case of $\operatorname{GL}_{2}$-type” abelian varieties, prove a result on the images of their Galois representations, and give a criterion, distinct from the definition, to identify their ordinary primes.
@@
which we call the $\ell$-adic Galois representation of AA. [^2] Due to the symplectic pairing on $V_{\ell}(A)$ coming from the polarization of $A$, the image of $\rho_{A,\ell}$ actually lies in $\operatorname{GSp}_{2g}(\mathbb{Q}_{\ell})$. This Galois representation satisfies that for all primes $v$ of $F$ with $v\nmid N_{A}\ell$, $\rho_{A,\ell}$ is unramified at $v$, and the characteristic polynomial
@@
    $A/\overline{F}$ is of type I or II in Albert’s classification and $D:=\operatorname{End}_{\overline{F}}(A)\otimes\mathbb{Q}=\operatorname{End}_{F}(A)\otimes\mathbb{Q}$.
@@
Let $A/F$ be an abelian variety of dimension $g$ over a number field $F$. Assume that $A$ is absolutely simple, $\operatorname{End}_{\overline{F}}(A)\otimes\mathbb{Q}$ contains a number field $K$ of degree gg, [^5] Note that this is weaker than the assumption $K\hookrightarrow\operatorname{End}_{F}(A)\otimes\mathbb{Q}$ in Theorem 1.3. and the center of $\operatorname{End}_{\overline{F}}(A)\otimes\mathbb{Q}$ is a totally real field $K^{\prime}$. Then, there exists a finite extension $F^{\prime}/F$ such that for all sufficiently large prime $\ell$, we have, up to conjugation, that
@@
If we replace $A$ by any abelian variety in its isogeny class over some finite extension of $F$, then the assumptions are still satisfied. Also, for all sufficiently large prime $\ell$, the image of the $\ell$-adic Galois representation remains unchanged over the field where the isogeny is defined. Since $A$ is isogenous over a finite extension of $F$ to a principally polarized abelian variety, we may assume that $A$ is principally polarized over $F$.
Now, let $F^{\prime}/F$ be a finite extension of number fields such that  [^6] The existence of this finite extension can be deduced from [Sil92, Theorem 4.1] and [SZ95, Lemma 2.7 and Remark 3.1] (see also [Ser13]).
@@
    $D:=\operatorname{End}_{\overline{F}}(A)\otimes\mathbb{Q}=\operatorname{End}_{F^{\prime}}(A)\otimes\mathbb{Q}$;
@@
By the assumptions and (2), $A/F^{\prime}$ is simple and of $\operatorname{GL}_{2}$-type, and by [Wu20, Proposition 1.7], $A/F^{\prime}$ is not of type III. As the center $K^{\prime}$ of $D$ is assumed to be totally real, by [Wu20, Proposition 1.7], we have either $D=K^{\prime}(=K)$ or $D$ is a division quaternion algebra over $K^{\prime}$ with $K^{\prime}:\mathbb{Q}]=g/2$. Putting all things together, it is then easy to check that $A/F^{\prime}$ is of class $\mathcal{A}$, so by [BGK06, Theorem 6.16 (6.18)], we have $\operatorname{SL}_{2}(\mathcal{O}_{K^{\prime}}\otimes\mathbb{Z}_{\ell})\subseteq\rho_{A,\ell}(\operatorname{Gal}(\overline{F^{\prime}}/F^{\prime}))$ for sufficiently large prime $\ell$. On the other hand, by the assumptions, $A/F^{\prime}$ is (geometrically) of the first kind. As the center $K^{\prime}$ of $D$ is totally real, by Proposition 3.4 (and gathering all primes $\lambda^{\prime}$ of $K^{\prime}$ with $\lambda^{\prime}|\ell$), we have $\operatorname{det}\rho_{A,\ell}|_{\operatorname{Gal}(\overline{F}/F^{\prime})}=\chi_{\ell}$ for all rational primes $\ell$.
@@
Let $A$ be an abelian variety of dimension $g$ and conductor $N_{A}$ defined over a number field $F$. We say that a prime $v$ of $F$ is an ordinary prime of $A/F$ if $v\nmid N_{A}$ and $p\nmid a_{v,g}(A)$, where $p$ is the rational prime lying below $v$ and $a_{v,g}(A)$ is the coefficient of the term $X^{g}$ of $P_{A,v}(X)$ defined by Equation 1.
@@
Since $\{\sigma:K\rightarrow\mathbb{C}\}=g$, it is easy to see that the coefficient $a_{v,g}(A)$ of the term $X^{g}$ is of the form
@@ 3.2. Galois representations of modular forms
such that for every rational prime $p\nmid\ell N$, $\rho_{f,\ell}$ is unramified at $p$ and
@@
By the Ramanujan conjecture, it is known that $|\iota(a_{p}(f))|_{\mathbb{C}}\leq 2p^{\frac{k-1}{2}}$ for all embeddings $\iota:K\hookrightarrow\mathbb{C}$.
@@
Since $\{\sigma:K_{f}\rightarrow\mathbb{C}\}=[K_{f}:\mathbb{Q}]=:g$, it is easy to see that the coefficient $a_{p,g}(A_{f})$ of the term $X^{g}$ is of the form
@@
where we write $K:\mathbb{Q}]=g$. By the assumptions on $v$ and $p$, we have $\mathrm{N}(v)=p$ and $\mathrm{N}(\lambda)=p^{g}$. Hence, $p^{g}|\mathrm{N}((a_{v}))$ with $\mathrm{N}((a_{v}))\leq 2^{g}p^{g/2}$. As $p\geq 5$, this implies that $\mathrm{N}((a_{v}))=0$ so $a_{v}=0$. ∎
@@
By Lemma 4.1, if $v\in S_{1}$, then $a_{v}=0$ or $v\in S^{\mathrm{ord}}$. Since $\{v\in\Sigma_{F}:v\notin S\text{ and }a_{v}=0\}$ is assumed to have density 0, $S_{1}$ thus has the same density as $S_{1}\cap\{v\in\Sigma_{F}:v\notin S\text{ and }a_{v}\neq 0\}\subseteq S^{\mathrm{ord}}$, so it suffices to prove that $S_{1}$ has positive density.
@@
$\#\{v\in\Sigma_{F}:\mathrm{N}(v)\leq X\}\sim\frac{X}{\log X}\sim\#\{p:p\leq X\}.$
@@
The proof of Theorem 1.7 follows a similar approach to that of Theorem 1.8. We first prove a lemma showing that under the assumption on the rational prime $p$, almost every prime $v$ of $F$ above $p$ is either an ordinary prime or belongs to a special set. In the case of Theorem 1.8, the special set is $\{v\in\Sigma_{F}:v\notin S\text{ and }a_{v}=0\}$, which has density 0 by assumptions. For Theorem 1.7, the special set is $S(c)$ (see Equation 4 for its precise definition) or, more accurately, a finite union of such sets. It is much trickier to show that the density of $S(c)$ is 0. Therefore, we first proceed with the proof assuming that $S(c)$ has density 0, and then, in Section 5.2, we discuss its density properly.
@@
where we write $K:\mathbb{Q}]=g$. By the assumptions on $v$ and $p$, we have $\mathrm{N}(v)=p$ and $\mathrm{N}(\lambda)=p^{g/2}$ ($g$ is necessarily even in this case). Hence, $p^{g/2}|\mathrm{N}((a_{v}))$ with $\mathrm{N}((a_{v}))\leq 2^{g}p^{g/2}$, implying that $\mathrm{N}_{K/\mathbb{Q}}(a_{v})^{2}=c\mathrm{N}(v)^{g}$ for some $c\in\mathbb{Z}$ with $0\leq c\leq 2^{2g}$. ∎
@@
Suppose for now that $S(c)$ has density 0, which will be proved in Corollary 5.7. We will show under this assumption that Theorem 1.7 holds. Let $T_{2}$ be the set of rational primes $p$ such that
@@
By Lemma 5.1, if $v\in S_{2}$, then $v\in S^{\mathrm{ord}}$ or $v\in S(c)$ for some $c\in\mathbb{Z}$ with $0\leq c\leq 2^{2g}$. As $S(c)$ and hence $\bigcup_{0\leq c\leq 2^{2g}}S(c)$ have density 0, it follows that $S_{2}$ has the same density as $S_{2}\cap S^{\mathrm{ord}}\subseteq S^{\mathrm{ord}}$, so it suffices to prove that $S_{2}$ has positive density. Now, the same argument as before gives that
@@ 5.2. Density of $S(c)$
The main goal of this subsection is to show that $S(c)$ has density 0 for each $c\in\mathbb{Z}$ with $0\leq c\leq 2^{2g}$, so that Theorem 1.7 holds by Section 5.1. To prove this, we need to use the surjective property of the family of Galois representations $(\rho_{\ell})_{\ell\in T}$.
Fix a rational prime $\ell$ that splits completely in $K$ and fix an embedding $\iota_{\ell}:\overline{\mathbb{Q}}\hookrightarrow\overline{\mathbb{Q}_{\ell}}$. Write $K^{\prime}:\mathbb{Q}]=g^{\prime}$ and let $\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{g^{\prime}}:K^{\prime}\hookrightarrow\overline{\mathbb{Q}}$ be all the embeddings of $K^{\prime}$ into $\overline{\mathbb{Q}}$. As $\ell$ splits completely in $K$ and hence in $K^{\prime}$, by Lemma 2.2, $\{\lambda^{\prime}_{i}:=\lambda^{\prime}_{\sigma^{\prime}_{i}}\mid 1\leq i\leq g^{\prime}\}$ is an enumeration of all the places $\lambda^{\prime}|\ell$ of $K^{\prime}$. In particular, we may simply make the identification $K^{\prime}_{\lambda^{\prime}_{i}}=\widehat{\iota_{\ell}\sigma^{\prime}_{i}(K^{\prime})}$$\mathbb{Q}_{\ell}\subseteq\overline{\mathbb{Q}_{\ell}}$ so that the natural map $K^{\prime}\rightarrow K^{\prime}_{\lambda^{\prime}_{i}}$ is given by $\alpha\mapsto\iota_{\ell}\sigma^{\prime}_{i}(\alpha)$. In this way,
@@
Write $K:K^{\prime}]=d$. Since $\ell$ splits completely in $K$, each $\lambda^{\prime}_{i}$ also splits completely in $K$. In particular, for every prime $\lambda|\lambda^{\prime}_{i}$ of $K$, we have $\mathcal{O}_{K^{\prime}_{\lambda^{\prime}_{i}}}=\mathcal{O}_{K_{\lambda}}$. Let us now enumerate the primes $\lambda|\ell$ of $K$ by $(\lambda_{ij})_{i=1,j=1}^{g^{\prime},\;d}$ such that $\prod_{j=1}^{d}\lambda_{ij}=\lambda^{\prime}_{i}$ and let $\sigma_{ij}:K\hookrightarrow\overline{\mathbb{Q}}$ be the embedding corresponding to the prime $\lambda_{ij}$ by Lemma 2.2. We also make the identification $K_{\lambda_{ij}}=\widehat{\iota_{\ell}\sigma_{ij}(K)}$$\mathbb{Q}_{\ell}\subseteq\overline{\mathbb{Q}_{\ell}}$ so that similarly
@@
$\overline{\rho^{\prime}_{\ell}}(\operatorname{Gal}(\overline{F}/F^{\prime}))=A_{\ell}(g^{\prime},d):=\big\{$ $(M_{ij})_{i=1,j=1}^{g^{\prime},\;d}\in\prod_{i=1}^{g^{\prime}}\prod_{j=1}^{d}\operatorname{GL}_{2}(\mathbb{F}_{\ell}):M_{ij}=M_{ij^{\prime}}\text{ for all }1\leq j,j^{\prime}\leq d,$
 $\hskip 19.91684pt\text{ and }\operatorname{det}M_{ij}=\operatorname{det}M_{i^{\prime}j^{\prime}}\text{ for all }1\leq i,i^{\prime}\leq g^{\prime}\text{ and }1\leq j,j^{\prime}\leq d\big\}.$
Now, since $F^{\prime}:F]$ is finite, there exists $n\in\mathbb{Z}^{+}$ such that $\tau^{n}\in\operatorname{Gal}(\overline{F}/F^{\prime})$ for all $\tau\in\operatorname{Gal}(\overline{F}/F)$. [^9] By considering all the cosets $\{\tau^{i}\operatorname{Gal}(\overline{F}/F^{\prime})\}_{i\in\mathbb{Z}}$, one can deduce that $\tau^{i}\in\operatorname{Gal}(\overline{F}/F^{\prime})$ for some $1\leq i\leq[F^{\prime}:F]$, so $n$ can be chosen as $F^{\prime}:F]!$. In particular, this means that
@@
 $\hskip 28.45274pt\text{ and }\operatorname{det}M_{ij}^{n}=\operatorname{det}M_{i^{\prime}j^{\prime}}^{n}\text{ for all }1\leq i,i^{\prime}\leq g^{\prime}\text{ and }1\leq j,j^{\prime}\leq d\big\}.$
@@
 $\hskip 14.22636pt\text{ and }\operatorname{det}M_{i1}^{n}=\operatorname{det}M_{i^{\prime}1}^{n}\text{ for all }1\leq i,i^{\prime}\leq d\text{ and }1\leq j\leq d\big\},$
@@
$C_{\ell}(g^{\prime},d,n,c):=\bigg\{$ $(M_{ij})_{i=1,j=1}^{g^{\prime},\;d}\in B_{\ell}(g^{\prime},d,n):\bigg(\prod_{i=1}^{g^{\prime}}\prod_{j=1}^{d}\operatorname{tr}M_{ij}\bigg)^{2}=\bar{c}\cdot\prod_{i=1}^{g^{\prime}}\prod_{j=1}^{d}\operatorname{det}M_{ij}\bigg\}.$
@@
We will now prove that the latter expression goes to 0 as $\ell\to\infty$, from which it follows that $S(c)$ has density 0.
@@
$\left\{N\in\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z}):\operatorname{det}N=d,\operatorname{tr}N=t\right\}=\ell^{2}+\ell\cdot\bigg(\frac{t^{2}-4d}{\ell}\bigg),$
@@
$|C_{\ell}(g^{\prime},d,n,c_{0})|\leq n^{g^{\prime}d}\sum_{c\in\mu_{n}(\mathbb{F}_{\ell})}|D_{\ell}(g^{\prime},d,n,c{\thinspace}\overline{c_{0}})|,$ (8)